OpenStudy (anonymous):

The coordinates of the vertices of two rectangles, Rectangle ABCD and Rectangle PQRS are given below. Rectangle ABCD – A(0, 8), B(8, 8), C(8, -4), D(0, -4) Rectangle PQRS – P(0, 4), Q(4, 4), R(4, -2), S(0, -2) Maria scaled both the rectangles about their centers to create two congruent rectangles A′B′C′D′ and P′Q′R′S′. By which factor did she most likely scale Rectangle ABCD and Rectangle PQRS?

OpenStudy (anonymous):

ANSWERS Rectangle ABCD by 1 by 2 and Rectangle PQRS by 2 to create congruent rectangles of dimensions 4 x 3 units Rectangle ABCD by 2 and Rectangle PQRS by 3 to create congruent rectangle of dimensions 3 x 2 units Rectangle ABCD by 1 by 4 and Rectangle PQRS by 4 to create congruent rectangles of dimensions 4 x 3 units Rectangle ABCD by 1 by 4 and Rectangle PQRS by 1 by 2 to create congruent rectangles of dimensions 2 x 3 units

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

start by centering both rectangles at the origin if they arent yet

OpenStudy (amistre64):

might start with a corner, then readjust

OpenStudy (amistre64):

lets corner them at the origin first Rectangle ABCD – A(0, 8), B(8, 8), C(8, -4), D(0, -4) -8 -8 -8 -8 --------------------------- 0 0 8 0 8 -12 0 -12 Rectangle PQRS – P(0, 4), Q(4, 4), R(4, -2), S(0, -2) -4 -4 -4 -4 -------------------------- 0 0 4 0 4 -6 0 -6

OpenStudy (amistre64):

now to center them 0 0 8 0 8 -12 0 -12 ^ ^ ^ ^ w = 8 so -4 h = -12 so +6 ------------------------------- -4 6 4 6 4 -6 -4 -6 0 0 4 0 4 -6 0 -6 ^ ^ ^ ^ w = 4 so -2 h = -6 so +3 --------------------------------- -2 3 2 3 2 -3 -2 -3

OpenStudy (amistre64):

now all we have to do is compare a set of similar corners: -4 6 to -2 3

OpenStudy (anonymous):

is there anything else?

OpenStudy (amistre64):

only thing left to do is to compare the options with these 2 points to see what fits ... im a little offput by the doubles: by 1 by 2; means what?

OpenStudy (anonymous):

is it the first answer????

OpenStudy (anonymous):

........or the last answer....

OpenStudy (amistre64):

i cant read the options .... they have double speak in them that is muddling things up

OpenStudy (anonymous):

Rectangle ABCD by 1 by 4 and Rectangle PQRS by 1 by 2 to create congruent rectangles of dimensions 2 x 3 units........is this the answer?

OpenStudy (amistre64):

what does: by 1 by 4 mean ?????

OpenStudy (amistre64):

... by 1/4 and ... by 1/2 maybe?

OpenStudy (anonymous):

idk here are the answers Rectangle ABCD by 1 by 2 and Rectangle PQRS by 2 to create congruent rectangles of dimensions 4 x 3 units Rectangle ABCD by 2 and Rectangle PQRS by 3 to create congruent rectangle of dimensions 3 x 2 units Rectangle ABCD by 1 by 4 and Rectangle PQRS by 4 to create congruent rectangles of dimensions 4 x 3 units Rectangle ABCD by 1 by 4 and Rectangle PQRS by 1 by 2 to create congruent rectangles of dimensions 2 x 3 units

OpenStudy (amistre64):

well, if you cant read the options, and i cant make sense of them ... then we are at a stand still

OpenStudy (amistre64):

reposting the same muddled jargon over and over again is NOT clarifying the issue ....

OpenStudy (amistre64):

the best i can determine is: n(4,6) = k(2,3 4n = 2k when n/k = 1/2

OpenStudy (amistre64):

the only thing that might make a lick of sense out ofthe options, would then be the last one ...

OpenStudy (anonymous):

ok thanks

OpenStudy (amistre64):

yep ...